# Fast and Accurate Collimator-Detector Response Compensation in High-Energy SPECT Imaging with 1D Convolutions and Rotations

### v1

@@ 4 Discussion
In this work, a 1D-R PSF model for high energy SPECT reconstruction was proposed, and demonstrated to be both efficacious and fast for use in high-energy reconstruction of 225Ac data. The model was created using publicly shared SPECTPSFToolbox component of the PyTomography project; t is hoped that the shared toolbox and corresponding tutorials will permit the nuclear medicine community to develop custom PSF models for new isotope/collimator configurations as research in novel RPTs and imaging techniques continues to grow.
As shown in Figure 3, the proposed 1D-R only yields computational benefits when image reconstruction is performed on GPU. If reconstruction is performed on CPU, then PSF modeling should be implemented with FFT-based 2D convolutions. FFT-based 2D convolutions are still, however, three times faster on GPU than on CPU. Furthermore, if a GPU is used then the 1D-R model can be used for further computational advantages. The fastest implementation of PSF modeling on GPU (proposed 1D-R) is approximately one order of magnitude faster than the fastest implementation of PSF modeling on CPU (FFT-based 2D convolutions).
While the proposed 1D-R model is not able to perfectly capture the features of the PSF at all radial distances in Figure 2, it yields near identical reconstructions in Figures 4 and 5, especially when compared to reconstruction using the Gaussian kernel. In the low count scenario of the phantom study, which approximately represents a clinical count rate, the percent difference in the three spheres between the 1D-R and 2D model were small compared to the variability between separate acquisitions. In the patient example, the variability of counts between the 2D and 1D-R models in the bone lesions was similar to the spheres in the phantom: this suggests that any differences in predicted counts between the 2D and 1D-R models in the patient example is also negligible compared to differences that would be observed between separate acquisitions. While the 1D-R model and 2D model yielded similar contrast, the Gaussian model had significantly reduced contrast with differences of $-51.3~{}\%$, $-23.3~{}\%$, and $-56.7~{}\%$ compared to the 2D model in the bone lesions. While scaling the magnitude of the image reconstructed using the Gaussian PSF could artificially increase the number of counts in each bone ROI, it would have no impact on these substantial differences in contrast.
In the phantom example, the radial path of the detector approached 7.27.27.2~{}7.2cm at its minimum and 27.227.227.2~{}27.2cm at its maximum. Since the radius of the cylinder was 101010~{}10cm, this suggests that the PSF model was able to reasonably function between $0-37~{}$cm. Future work may seek to test the efficacy of the proposed 1D-R model for larger radial distances, as this may be required for larger patients.
While the functional form of Equations 9, 12, 15 were selected because they produced a reasonable approximation of the PSF, it may be the case that they can be improved upon. Use of rotation operations with the component $K_{B}^{(x;d;b)}K_{B}^{(y;d;b)}$ may permit more radial symmetry in the SSRF. Substitution of $K_{G}^{(x;d;b)}K_{G}^{(y;d;b)}$ with a small 2D kernel that represents the true aperture function of the GRF may remedy the under prediction in the center of the PSFs at large distances. Close inspection of the SIMIND data at large distances in Figure 2 reveals additional dim tails; these could be modeled by including another independent tail component in the model. A trade-off for adding additional features to the model, however, is an increase in computation time. The SPECTPSFToolbox Python library released along with this paper contains tutorials demonstrating how to create custom and fit parameterized PSF models. Users can obtain point source data using Monte Carlo programs (such as GATE [40] and SIMIND) or by using real scanners, and fit corresponding models to the acquired data. The models can then be imported to PyTomography for customized and scanner specific SPECT reconstruction. Users can independently evaluate the trade-off between adding model features, impact on reconstructed images, and increase in computation time.
@@
As discussed in Section 2.1.2, the kernel sizes were fixed to $N-1\times N-1$ for all $d$. For small $d$, where the PSF is also smaller, a kernel of this size might be excessive and could unnecessarily increase the computational time required. Future research exploring the reduction of computational time might experiment with using a source-detector distance dependent kernel size $N(d)$ for that matches the size of the PSF; this would further reduce the computational time required.
@@
Since reconstruction using the proposed 1D-R model more than halves reconstruction times compared to the 2D model, it may be preferable for reconstruction of 225Ac data. Both the 1D-R and 2D models, however, are only applicable when Equation 5 is used for SPECT system matrix modeling; this equation relies on the assumption that the attenuation probabilities vary little across the PSF. Due to the large spatial extent of the PSFs shown here, this assumption may be invalid, and Equation 4 may instead be required for accurate reconstruction of 225Ac data. This assumption largely depends on the density of the object being scanned. If the object had a region of high density which only rays on the outer edges of the PSF passed through, then the equal attenuation path assumption would be invalid. For patients and phantoms, where the density is roughly constant throughout the field of view, application of Equation 5 may be permitted. Meanwhile, the study of different density configurations and the applicability of Equation 5 remains to be investigated in future work.
@@
In conclusion, a fast and accurate implementation of high energy CDR modeling in SPECT imaging that uses 1D convolutions and rotations was developed, and tested on 225Ac reconstructions. The technique was implemented using the open source reconstruction library PyTomography, and was shown to speed up reconstruction times by more than a factor of two compared to conventional, 2D convolution based methods. Furthermore, the 1D-R method yielded near identical results to the 2D method, and the small differences were insignificant compared to the differences between reconstructions of separate noise realizations. The SPECTPSFToolbox python library was developed and publicly shared to permit others in the nuclear medicine community to develop custom isotope/collimator PSF models for use in the open-source reconstruction library PyTomography.

### v2

@@ 4 Discussion
This work proposes a 1DR-MC PSF model for high-energy SPECT reconstruction that demonstrates consistency with 2D-MC models, while also increasing computational efficiency for GPU-based reconstruction of 177Lu, 131I, and 225Ac at typical SPECT matrix sizes. The model was created using the publicly shared SPECTPSFToolbox component of the PyTomography project; it is hoped that the shared toolbox and corresponding tutorials will permit the nuclear medicine community to develop custom PSF models for new isotope/collimator configurations as research in novel RPTs and imaging techniques continues to grow.
@@
As shown in Figure 3, the proposed 1DR-MC only yields computational benefits when image reconstruction is performed on GPU. If reconstruction is performed on CPU, then PSF modeling is faster when implemented with FFT-based 2D convolutions. As discussed in Section 2.1.2, the kernel sizes were fixed to $N-1\times N-1$ for all $d$. For small $d$, where the PSF is also smaller, a kernel of this size might be excessive and could unnecessarily increase the computational time required. Future research exploring the reduction of computational time might experiment with using a source-detector distance dependent kernel size $N(d)$ for that matches the size of the PSF; this would further reduce the computational time required.
While the proposed 1DR-MC model is not able to perfectly capture all the PSF features in Figure 2, it yields qualitatively similar reconstructions of the MC data in Figure 4. Furthermore, as shown in Table 2, the RCs and contrast of the 1DR-MC and 2D-MC models are similar and always significantly greater than the 2DG model. This demonstrates the superior quantitative performance of the MC-based models, and justification for using the 1DR-MC over the 2D-MC method due to faster reconstruction times.
For the physical 225Ac phantom and patient use cases shown in Figures 5 and 6, the 2D-MC and 1DR-MC models yield similar count quantification, especially when compared to the 2DG model. The quality metrics for 200 iterative updates shown in Table 2 further highlight this consistency. In the physical phantom high count case, use of the 1DR-MC method compared to the 2DG method increased the mean counts the spheres by factors of 2.06, 2.14, and 1.92 for the largest to smallest sphere respectively; this is somewhat similar to the MC study where the RCs increased by 2.17, 2.45, and 2.11 for spheres of the same size. While calibration factors were not available for the physical data, inference from the MC results suggest that the RCs of the physical data were significantly improved.
In the low count scenario of the physical phantom study, which approximately represents a clinical count rate, the percent difference in the three spheres between the 2D-MC and 1DR-MC models were small compared to the variability between the thirty-four separate acquisitions. This highlights that any systematic errors between use of the 2D-MC and 1DR-MC model are negligible compared to the precision of the imaging protocol, and thus that the 1DR-MC model may be suitable for standard 225Ac image reconstruction. In the patient example, it was similarly found that (i) the differences between the 2D-MC and 1DR-MC models in the bone lesions were small and (ii) the 2DG model had significantly reduced contrast relative to the 2D-MC and 1DR-MC models.
The functional form of Equations 9, 12, and 15 were selected because they produced a reasonable approximation of the PSF, but they can be further developed. Use of rotation operations with the component $K_{B}^{(x;d;b)}K_{B}^{(y;d;b)}$ may permit more radial symmetry in the SSRF. Substitution of $K_{G}^{(x;d;b)}K_{G}^{(y;d;b)}$ with a small 2D kernel that represents the true aperture function of the GRF may also improve the accuracy of the PSF model. The 177Lu PSF data in Figure 2 reveals additional dim tails; these could be modeled by including another independent tail component in the model. A trade-off for adding additional features to the model, however, is an increase in computation time. The SPECTPSFToolbox Python library released along with this paper contains tutorials demonstrating how to create custom and fit parameterized PSF models. Users can obtain point source data using Monte Carlo programs (such as GATE [41] and SIMIND) or by using real scanners, and fit corresponding models to the acquired data. The models can then be imported to PyTomography for customized and scanner specific SPECT reconstruction. Users can independently evaluate the trade-off between adding model features, impact on reconstructed images, and increase in computation time.
@@
Since GPU-accelerated reconstruction using the proposed 1DR-MC model reduces reconstruction times by more than a factor of two compared to the 2D-MC model, it may be preferable for reconstruction of 225Ac data for typical SPECT matrix sizes. Both the 1DR-MC and 2D-MC models, however, are only applicable when Equation 5 is used for SPECT system matrix modeling; this equation relies on the assumption that the attenuation probabilities vary little across the PSF. Due to the large spatial extent of the PSFs shown here, this assumption may be invalid, and Equation 4 may instead be required for accurate reconstruction. This assumption largely depends on the density of the object being scanned. If the object had a region of high density which only rays on the outer edges of the PSF passed through, then the equal attenuation path assumption would be invalid. For patients and phantoms, where the density is roughly constant throughout the field of view, application of Equation 5 may be permitted. Meanwhile, the study of different density configurations and the applicability of Equation 5 remains to be investigated in future work.
5 Conclusion
@@
In conclusion, a computationally efficient implementation of high energy CDR modeling in SPECT imaging was developed and tested on 177Lu, 131I, and 225Ac reconstructions. The technique was implemented using the open source reconstruction library PyTomography, and was shown to speed up reconstruction times by more than a factor of two compared to conventional, 2D convolution based methods ($272.0$ s vs. $704.0$ s for patient data). Furthermore, the proposed 1DR-MC method yielded near identical results to the reference 2D-MC approach, and the small differences for 225Ac were insignificant compared to the differences between reconstructions of separate noise realizations. The SPECTPSFToolbox python library was developed and publicly shared to permit others in the nuclear medicine community to develop custom isotope/collimator PSF models for use in the open-source reconstruction library PyTomography.